OpenStudy (anonymous):

Determine the measures of the unknown angles in the illustrations below...give me a minute to draw it out. It's in degrees, and that is a triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):

it's n degrees.

OpenStudy (anonymous):

53

OpenStudy (anonymous):

lol i need to know how to do it. it would be really helpful

OpenStudy (anonymous):

See the sum of the three interior angles of a triangle is ALWAYS 180 degree. You have been given measures of two angles and let the third angle be X then X + 82 + 45 = 180 X +127 = 180 X = 180 - 127 = 53 degree

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

thanks